A stoichiometric mixture of ethanol (C2H6O), octane (C8H18), and air initially at T1 = 300 K, P1 = 100 kPa, P2 =4500 kPa, and V1 = 5 liters (0.005 m3) is compressed in a piston-cylinder assembly in a polytropic process where n = 1.2

What is the final volume V2 (m3 ), the initial total volume of moles (kmol), and the final Temperature T2?

NOTE: Universal gas constant: Ru = 8.314 k]/ kmolK

Please show ALL steps and units with unit conversions.
Answer #1

For a polytopic process, PVn = constant or P1V1n = P2V2n ...... eqn 1

and (P1/P2)n-1/n = T1/T2 ......... eqn 2

From equation 2,

(100/4500)0.2/1.2 = 300/T2

So, T2 = 566K

From equation 1,

100 × 0.0051.2 = 4500 × V21.2

or, V21.2 = 3.85 * 10-5

or, 1.2logV2 = 3.85 * 10-5

or, V2 = 2.1 * 10-4 m3

We know, n = P1V1/ RT1

= 0.987 atm * 5 lit / (0.082 lit.atm.K-1.mol-1) . 300K ...... [100kPa = 0.98 atm]

= 0.2 mol